# Density estimation via mixture discrepancy and moments

### v1

@@ 2 The DSP-mix method
Compared with the star discrepancy adopted in DSP, the mixture discrepancy used in DSP-mix has the following advantages [11].
  • 1.
@@
    and
    $D^{\text{mix}}(\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}$ (10)
    $=$ $\left(\frac{19}{12}\right)^{d}-\frac{2}{n}\sum_{i=1}^{n}\left(\frac{5}{3}-\frac{1}{4}\left|\frac{1}{2}-y_{i,j_{0}}\right|-\frac{1}{4}\left|\frac{1}{2}-y_{i,j_{0}}\right|^{2}\right)\prod_{j\neq j_{0}}\left(\frac{5}{3}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|^{2}\right)$
    $+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}\left(\frac{15}{8}-\frac{1}{4}\left|\frac{1}{2}-y_{i,j_{0}}\right|-\frac{1}{4}\left|\frac{1}{2}-y_{k,j_{0}}\right|-\frac{3}{4}\left|y_{i,j_{0}}-y_{k,j_{0}}\right|+\frac{1}{2}\left|y_{i,j_{0}}-y_{k,j_{0}}\right|^{2}\right)$
    $\prod_{j\neq j_{0}}\left(\frac{15}{8}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{kj}-\frac{1}{2}\right|-\frac{3}{4}\left|y_{ij}-y_{kj}\right|+\frac{1}{2}\left|y_{ij}-y_{kj}\right|^{2}\right)$
    $=$ $D^{\text{mix}}(\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}.$
    For the two point sets in Figures 1(a) and 1(b), their mixture discrepancies are both $\frac{2719}{18432}$;
  • 3.
@@
    $\widetilde{y}_{ij}=\begin{cases}y_{ij},&j\neq i_{0}\,\text{and}\,j\neq j_{0},\\ \cos\left(-\pi/2\right)(y_{i,i_{0}}-0.5)-\sin\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=y_{i,j_{0}},&j=i_{0},\\ \sin\left(-\pi/2\right)(y_{i,i_{0}}-0.5)+\cos\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=1-y_{i,i_{0}},&j=j_{0},\end{cases}$
    and it can be easily proved that $D^{\text{mix}}(\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}=D^{\text{mix}}(\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$, analogous to Eq.(10). For the two point sets in Figures 2(a) and 2(b), their mixture discrepancies are both $\frac{301}{1024}$.
3 The MSP method
MSP still adopts the framework of Algorithm 1. However, instead of using the discrepancy as the uniformity measurement, it judges whether the particle set $S_{l}$ in $\Omega_{l}$ is close to a uniform distribution by comparing whether the moments of the discrete point measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{y}_{i}}$ and the uniform measure are near enough. Suppose that the $d$-dimensional random variable $\bm{y}=(y_{1},y_{2},\dots,y_{d})$ follows the probability density $q(\bm{y})$, the $d$-dimensional index $\boldsymbol{k}=(k_{1},k_{2},\dots,k_{d})$, $\bm{y}^{\boldsymbol{k}}=y_{1}^{k_{1}}y_{2}^{k_{2}}\dots y_{d}^{k_{d}}$, $\mathbb{N}_{0}=\{0,1,2,\dots\}$, then the moment of order $|\boldsymbol{k}|=\sum\limits_{i=1}^{d}k_{i}$ of the probability density $q(\bm{y})$ is defined as
@@ 3 The MSP method
For two probability densities defined on a compact set, we can judge whether they are close by comparing their moment sequences. In actual numerical implementation, we discover that conducting comparisons up to the order of $|\boldsymbol{k}|\leq 2$ is capable of attaining satisfactory numerical effects, that is, only the expectation and the covariance matrix are considered. Denote the expectation of the uniform measure on the sub-domain $\Omega_{l}$ as $\boldsymbol{\mu}^{(l)}=(\mu^{(l)}_{1},\dots,\mu^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix as $\boldsymbol{\Sigma}^{(l)}=(\Sigma_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. The expectation of the particle subset $S_{l}$ is $\hat{\boldsymbol{\mu}}^{(l)}=(\hat{\mu}^{(l)}_{1},\dots,\hat{\mu}^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix is $\boldsymbol{\hat{\Sigma}}^{(l)}=(\hat{\Sigma}_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. They have the explicit calculation expressions,
@@
$\Sigma_{jj}^{(l)}$ $=\frac{\left(b_{j}^{(l)}-a_{j}^{(l)}\right)^{2}}{12},\quad j=1,\dots,d,$
$\Sigma_{ij}^{(l)}$ $=0,\quad i\neq j,$
$\hat{\boldsymbol{\mu}}^{(l)}$ $=\frac{1}{n}\sum_{i=1}^{n}\bm{y}_{i}^{(l)},$
$\boldsymbol{\hat{\Sigma}}^{(l)}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)^{T}.$
@@
$\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|<\epsilon_{1}\left(b_{j}^{(l)}-a_{j}^{(l)}\right),\quad j=1,\dots,d,$ (13)
$\left|\Sigma_{jj}^{(l)}-\hat{\Sigma}_{jj}^{(l)}\right|<\epsilon_{2}\left|\Sigma_{jj}^{(l)}\right|,\quad j=1,\dots,d,$
$\left|\hat{\Sigma}_{ij}^{(l)}\right|<\epsilon_{3},\quad i\neq j,$
where $\epsilon_{1},\epsilon_{2},\epsilon_{3}>0$ are tolerance parameters. If Eq. (13) is satisfied, MSP terminates the sequential partition; Otherwise we consider that $S_{l}$ cannot be approximated as the uniform distribution in Line 7 of Algorithm 1. Obviously, MSP has the following advantages over DSP.
  • 1.
@@
    $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\widetilde{\hat{\Sigma}}_{j_{0}j_{0}}^{(l)}\right|$ (16)
    $=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{i=1}^{n}\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}})\right)^{2}\right|$
    $=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{k=1}^{n}\left(y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}y_{k,j_{0}}\right)^{2}\right|$
    $=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\hat{\Sigma}_{j_{0}j_{0}}^{(l)}\right|.$
@@
    $\left|\widetilde{\hat{\Sigma}}_{ij_{0}}^{(l)}\right|$ (17)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{h,j_{0}})\right)\right|$
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\right|$
    $=$ $\left|\hat{\Sigma}_{ij}^{(l)}\right|.$
    The situation with $i=j_{0}$ and $j\neq j_{0}$ is analogous. Combining Eqs. (15), (16) and (17) together, the reflection transformation has no impact on the termination criterion (13). For Figures 1(a) and 1(b), the differences between the expectations and covariance matrices of the two groups of particles and those of the uniform measure are
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(\frac{1}{12},\frac{1}{12}\right),$ (18)
    $\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{48}&\frac{1}{48}\\ \frac{1}{48}&\frac{1}{48}\end{pmatrix}.$
@@
    $\widetilde{y}_{ii_{0}}$
    $=$ $\cos\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)-\sin\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}$
@@
    $\widetilde{y}_{ij_{0}}$
    $=$ $\sin\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)+\cos\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}$
@@
    $\left|\widetilde{\hat{\Sigma}}_{i_{0}j_{0}}^{(l)}\right|$ (21)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{h,i_{0}})\right)\right|$
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i_{0}}\right)\right|$
    $=$ $\left|\hat{\Sigma}_{i_{0}j_{0}}^{(l)}\right|.$
@@
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0.25,0\right),$ (22)
    $\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{5}{48}&0\\ 0&\frac{1}{24}\end{pmatrix}.$
@@
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0,0.25\right),$ (23)
    $\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{24}&0\\ 0&\frac{5}{48}\end{pmatrix}.$
    Obviously, Eqs. (22) and (23) merely involve a rearrangement of elements and thus does not affect the termination criterion (13). That is, the origin $\boldsymbol{0}$ in MSP is indistinctive, and there will be no such unreasonable phenomena of the star discrepancy in Figures 1 and 2.

### v2

@@ 2.3 Relaxation by the mixed discrepancy
Compared with the star discrepancy adopted in DSP, the mixture discrepancy has the following advantages [26].
  1. (1)
@@
    and $\left(D^{\text{mix}}(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}=\left(D^{\text{mix}}(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$. For the two point sets in Figures 2(a) and 2(b), their mixture discrepancies are both $\frac{2719}{18432}$;
  3. (3)
@@
    $\widetilde{y}_{ij}=\begin{cases}y_{ij},&j\neq i_{0},j_{0},\\ \cos\left(-\pi/2\right)(y_{i,i_{0}}-0.5)-\sin\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=y_{i,j_{0}},&j=i_{0},\\ \sin\left(-\pi/2\right)(y_{i,i_{0}}-0.5)+\cos\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=1-y_{i,i_{0}},&j=j_{0},\end{cases}$
    and it can be easily proved that $(D^{\text{mix}}\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}=(D^{\text{mix}}\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$. For the two point sets in Figures 3(a) and 3(b), their mixture discrepancies are both $\frac{301}{1024}$.
3 MSP: Moment-based Sequential Partition
MSP still adopts the framework of Algorithm 1. Instead of using the discrepancy as the uniformity measurement, MSP judges whether the particle set $S_{l}$ in $\Omega_{l}$ is close to a uniform distribution by comparing whether the moments of the discrete point measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{y}_{i}}$ and the uniform measure are close enough.
The motivation comes from the following fact: Suppose that the $d$-dimensional random variable $\bm{y}=(y_{1},y_{2},\dots,y_{d})$ follows the probability density $q(\bm{y})$, the $d$-dimensional index $\boldsymbol{k}=(k_{1},k_{2},\dots,k_{d})$, $\bm{y}^{\boldsymbol{k}}=y_{1}^{k_{1}}y_{2}^{k_{2}}\dots y_{d}^{k_{d}}$, $\mathbb{N}_{0}=\{0,1,2,\dots\}$, then the moment of order $|\boldsymbol{k}|=\sum\limits_{i=1}^{d}k_{i}$ of the probability density $q(\bm{y})$ is defined as
@@ 3 MSP: Moment-based Sequential Partition
As a consequence, for two probability densities defined on a compact set, we can judge whether they are close by comparing their moment sequences. Detailedly, for the test function $f(\bm{y})$,
$\begin{split}&\int_{\Omega}f(\mathbf{y})\hat{p}(\mathbf{y})\mathrm{d}\mathbf{y}-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\\ =&\int_{\Omega}f(\mathbf{y})\hat{p}(\mathbf{y})\mathrm{d}\mathbf{y}-\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})+\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\\ \leq&\sum_{l=1}^{L}\frac{n_{l}}{N}\Big|\frac{1}{\textup{vol}(\Omega_{l})}\int_{\Omega_{l}}f(\mathbf{y})\mathrm{d}\mathbf{y}-\frac{1}{n_{l}}\sum_{\mathbf{y}_{i}^{(l)}\in S_{l}}f(\mathbf{y}_{i}^{(l)})\Big|\\ &+\Big|\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\Big|.\end{split}$ (13)
Thus we can choose $f(\bm{y})=\bm{y}^{k}$ to control the moment sequence in the density estimator, while the second term depends only on the Monte Carlo sampling error.
In actual numerical implementation, we discover that conducting comparisons up to the order of $|\boldsymbol{k}|\leq 2$ is capable of attaining satisfactory numerical effects, that is, only the expectation and the covariance matrix are considered. Denote the expectation of the uniform measure on the sub-domain $\Omega_{l}$ as $\boldsymbol{\mu}^{(l)}=(\mu^{(l)}_{1},\dots,\mu^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix as $\boldsymbol{\Sigma}^{(l)}=(\Sigma_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. The expectation of the particle subset $S_{l}$ is $\hat{\boldsymbol{\mu}}^{(l)}=(\hat{\mu}^{(l)}_{1},\dots,\hat{\mu}^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix is $\boldsymbol{\hat{\Sigma}}^{(l)}=(\hat{\Sigma}_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. They have the explicit calculation expressions,
@@
$\Sigma_{jj}^{(l)}$ $=\frac{\left(b_{j}^{(l)}-a_{j}^{(l)}\right)^{2}}{12},\quad j=1,\dots,d,\quad\textup{and}\quad\Sigma_{ij}^{(l)}$ $=0,\quad i\neq j,$
and
$\hat{\boldsymbol{\mu}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\bm{y}_{i}^{(l)},\quad\boldsymbol{\hat{\Sigma}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)^{T}.$ (15)
@@
$\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|<\theta\left(b_{j}^{(l)}-a_{j}^{(l)}\right),\quad j=1,\dots,d,$ (16)
$\left|\Sigma_{jj}^{(l)}-\hat{\Sigma}_{jj}^{(l)}\right|<\theta\left|\Sigma_{jj}^{(l)}\right|,\quad j=1,\dots,d,\quad\textup{and}\quad\left|\hat{\Sigma}_{ij}^{(l)}\right|<\theta,\quad i\neq j,$
where we choose the same tolerance parameter $\theta$ in Eq. (7) used by both DSP and DSP-mix for the sake of comparison. If Eq. (16) is satisfied, MSP terminates the sequential partition. Otherwise we consider that $S_{l}$ cannot be approximated as the uniform distribution in Line 7 of Algorithm 1. Obviously, MSP has the following advantages over DSP.
  1. (1)
@@
    $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\widetilde{\hat{\Sigma}}_{j_{0}j_{0}}^{(l)}\right|=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{i=1}^{n}\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}})\right)^{2}\right|$ (19)
    $=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{k=1}^{n}\left(y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}y_{k,j_{0}}\right)^{2}\right|=\left|\Sigma_{j_{0}j_{0}}^{(l)}-\hat{\Sigma}_{j_{0}j_{0}}^{(l)}\right|.$
@@
    $\left|\widetilde{\hat{\Sigma}}_{ij_{0}}^{(l)}\right|=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{h,j_{0}})\right)\right|$ (20)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\right|=\left|\hat{\Sigma}_{ij}^{(l)}\right|.$
    The situation with $i=j_{0}$ and $j\neq j_{0}$ is analogous. Combining Eqs. (18), (19) and (20) together, the reflection transformation has no impact on the termination criterion (16). For Figures 2(a) and 2(b), the differences between the expectations and covariance matrices of the two groups of particles and those of the uniform measure are
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(\frac{1}{12},\frac{1}{12}\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{48}&\frac{1}{48}\\ \frac{1}{48}&\frac{1}{48}\end{pmatrix}$, (21)
@@
    $\widetilde{y}_{ii_{0}}=$ $\cos\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)-\sin\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}$
@@
    $\widetilde{y}_{ij_{0}}=$ $\sin\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)+\cos\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}$
@@
    $\left|\widetilde{\hat{\Sigma}}_{i_{0}j_{0}}^{(l)}\right|=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{h,i_{0}})\right)\right|$ (24)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i_{0}}\right)\right|=\left|\hat{\Sigma}_{i_{0}j_{0}}^{(l)}\right|.$
@@
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0.25,0\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{5}{48}&0\\ 0&\frac{1}{24}\end{pmatrix}$. (25)
@@
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0,0.25\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{24}&0\\ 0&\frac{5}{48}\end{pmatrix}$. (26)
    Obviously, Eqs. (25) and (26) merely involve a rearrangement of elements and thus does not affect the termination criterion (16). That is, the origin $\boldsymbol{0}$ in MSP is indistinctive, and there will be no such unreasonable phenomena of the star discrepancy in Figures 2 and 3.